\begin{thebibliography}{10}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Agarwal et~al.(2021)Agarwal, Chakraborti, Fu, Gros, Lin, Maene,
  Talamadupula, Teng, and White}]{nlc2cmd}
Mayank Agarwal, Tathagata Chakraborti, Quchen Fu, David Gros, Xi~Victoria Lin,
  Jaron Maene, Kartik Talamadupula, Zhongwei Teng, and Jules White. 2021.
\newblock \href {https://proceedings.mlr.press/v133/agarwal21b.html} {Neurips
  2020 nlc2cmd competition: Translating natural language to bash commands}.
\newblock In \emph{Proceedings of the NeurIPS 2020 Competition and
  Demonstration Track}, volume 133 of \emph{Proceedings of Machine Learning
  Research}, pages 302--324. PMLR.

\bibitem[{Luong et~al.(2015)Luong, Pham, and
  Manning}]{luong-etal-2015-effective}
Thang Luong, Hieu Pham, and Christopher~D. Manning. 2015.
\newblock \href {https://doi.org/10.18653/v1/D15-1166} {Effective approaches to
  attention-based neural machine translation}.
\newblock In \emph{Proceedings of the 2015 Conference on Empirical Methods in
  Natural Language Processing}, pages 1412--1421, Lisbon, Portugal. Association
  for Computational Linguistics.

\bibitem[{Papadimitriou and
  Jurafsky(2020)}]{papadimitriou-jurafsky-2020-learning}
Isabel Papadimitriou and Dan Jurafsky. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.554} {{L}earning
  {M}usic {H}elps {Y}ou {R}ead: {U}sing transfer to study linguistic structure
  in language models}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 6829--6839, Online. Association
  for Computational Linguistics.

\bibitem[{Shi et~al.(2021)Shi, Ng, Wang, Zhu, Li, Wang, Nogueira~dos Santos,
  and Xiang}]{gap}
Peng Shi, Patrick Ng, Zhiguo Wang, Henghui Zhu, Alexander~Hanbo Li, Jun Wang,
  Cicero Nogueira~dos Santos, and Bing Xiang. 2021.
\newblock \href {https://ojs.aaai.org/index.php/AAAI/article/view/17627}
  {Learning contextual representations for semantic parsing with
  generation-augmented pre-training}.
\newblock \emph{Proceedings of the AAAI Conference on Artificial Intelligence},
  35(15):13806--13814.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{vaswani2017}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, \L~ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock \href
  {https://proceedings.neurips.cc/paper/2017/file/3f5ee243547dee91fbd053c1c4a845aa-Paper.pdf}
  {Attention is all you need}.
\newblock In \emph{Advances in Neural Information Processing Systems},
  volume~30. Curran Associates, Inc.

\bibitem[{Xu et~al.(2020)Xu, Jiang, Yin, Vasilescu, and
  Neubig}]{xu-etal-2020-incorporating}
Frank~F. Xu, Zhengbao Jiang, Pengcheng Yin, Bogdan Vasilescu, and Graham
  Neubig. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.538} {Incorporating
  external knowledge through pre-training for natural language to code
  generation}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 6045--6052, Online. Association for
  Computational Linguistics.

\bibitem[{Yin et~al.(2018)Yin, Deng, Chen, Vasilescu, and Neubig}]{conala}
Pengcheng Yin, Bowen Deng, Edgar Chen, Bogdan Vasilescu, and Graham Neubig.
  2018.
\newblock \href {https://doi.org/https://doi.org/10.1145/3196398.3196408}
  {Learning to mine aligned code and natural language pairs from stack
  overflow}.
\newblock In \emph{International Conference on Mining Software Repositories},
  MSR, pages 476--486. ACM.

\bibitem[{Yu et~al.(2021{\natexlab{a}})Yu, Wu, Lin, bailin wang, Tan, Yang,
  Radev, richard socher, and Xiong}]{yu2021grappa}
Tao Yu, Chien-Sheng Wu, Xi~Victoria Lin, bailin wang, Yi~Chern Tan, Xinyi Yang,
  Dragomir Radev, richard socher, and Caiming Xiong. 2021{\natexlab{a}}.
\newblock \href {https://openreview.net/forum?id=kyaIeYj4zZ} {Gra{PP}a:
  Grammar-augmented pre-training for table semantic parsing}.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Yu et~al.(2021{\natexlab{b}})Yu, Zhang, Polozov, Meek, and
  Awadallah}]{yu2021score}
Tao Yu, Rui Zhang, Alex Polozov, Christopher Meek, and Ahmed~Hassan Awadallah.
  2021{\natexlab{b}}.
\newblock \href {https://openreview.net/forum?id=oyZxhRI2RiE} {{SC}ore:
  Pre-training for context representation in conversational semantic parsing}.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Yu et~al.(2018)Yu, Zhang, Yang, Yasunaga, Wang, Li, Ma, Li, Yao,
  Roman, Zhang, and Radev}]{spider}
Tao Yu, Rui Zhang, Kai Yang, Michihiro Yasunaga, Dongxu Wang, Zifan Li, James
  Ma, Irene Li, Qingning Yao, Shanelle Roman, Zilin Zhang, and Dragomir Radev.
  2018.
\newblock Spider: A large-scale human-labeled dataset for complex and
  cross-domain semantic parsing and text-to-sql task.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, Brussels, Belgium. Association for
  Computational Linguistics.

\end{thebibliography}

\pdfoutput=1
\documentclass[11pt]{article}

\usepackage{acl}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}

\usepackage{microtype}

\title{Structural Transfer Learning in NL-to-Bash Semantic Parsers}

\author{Kyle Duffy \\
  University of Oxford \\
  \texttt{kylejduffy3@gmail.com} \\ \And
  Satwik Bhattamishra \\
  University of Oxford \\
  \texttt{satwik.bmishra@stx.ox.ac.uk} \\ \AND
  Phil Blunsom \\
  University of Oxford \\
  \texttt{phil.blunsom@cs.ox.ac.uk} \\}

\usepackage{amsthm, amssymb, amsmath, bbm}

\usepackage{enumitem}
\newlist{checklist}{itemize}{2}
\setlist[checklist]{label=$\square$}
\usepackage{pifont}

\newenvironment{itemize*}%
	{\begin{itemize}%
		\setlength{\itemsep}{0pt}%
		\setlength{\parskip}{0pt}%
		\setlength{\parsep}{0pt}}%
	{\end{itemize}}
\newenvironment{enumerate*}%
	{\begin{enumerate}%
		\setlength{\itemsep}{0pt}%
		\setlength{\parskip}{0pt}%
		\setlength{\parsep}{0pt}}%
	{\end{enumerate}}
\newenvironment{checklist*}%
	{\begin{checklist}%
		\setlength{\itemsep}{0pt}%
		\setlength{\parskip}{0pt}%
		\setlength{\parsep}{0pt}}%
	{\end{checklist}}
\DeclareMathOperator{\ppl}{ppl}

\usepackage{lipsum}
\usepackage{booktabs}

\begin{document}
\maketitle

\section{Introduction}

Semantic parsing is the task of extracting semantic meaning from natural language. The meaning is parsed into a machine-readable format, which is typically a programming language. In this paper, we consider a semantic parsing system for the natural language to Bash (NLBash) task.

Modifying pre-training data with synthesized samples or external information can improve performance of semantic parsers; synthesizing pre-training data with synchronous context-free grammars has improved performance on the natural language to SQL (NLSQL) task \citep{yu2021grappa,yu2021score,gap}, and incorporating context from external web data has proven effective for the natural language to Python (NLPython) task \citep{xu-etal-2020-incorporating}.

However these systems are often language-specific or rule-based, and end up being difficult to scale, imprecisely targeted, and monolingual.
A deeper understanding of structural transfer to semantic parsing could suggest which datasets should be included during pre-training and broaden the scope of pre-training data to include multilingual or out-of-task samples.

We measure transferability to NLBash from 6 upstream tasks by computing a heuristic for semantic accuracy of Bash programs introduced by \citet{nlc2cmd}, and also accuracy and perplexity. We find that NLSQL transfers strongly to NLBash, NLPython has little marginal improvement over lexical mapping, and more data with longer upstream training does not always lead to more semantically accurate programs.

\section{Method}\label{sec:method}

We measure how well a transformer can model NLBash using weights learned from a different task. To do so, we follow \citet{papadimitriou-jurafsky-2020-learning} and freeze transformer weights after pre-training and then fine-tune a linear embeddings layer on NLBash.

Formally, an upstream source vocabulary $S_U$, upstream target vocabulary $T_U$, downstream source vocabulary $S_D$, and downstream target vocabulary $T_D$ are given along with their corresponding sets of sentences, which we denote as $S_U^*$, $T_U^*$, $S_D^*$, $T_D^*$.

A transformer model $Y$ of a fixed depth and hidden dimension is chosen and initialized. An upstream task $U$ and a downstream task $D$ are selected and given as a set of sentence pairings, $U\subseteq S_U^*\times T_U^*$, $D\subseteq S_D^*\times T_D^*$. The downstream task is split into a training set $D_{\text{train}}$, and a test set $D_\text{test}$.

An embedding layer $e_U = (e_U^s, e_U^t)$ is prepended to the transformer model $Y$, and the model is trained end-to-end on the upstream task to convergence. This is called the \emph{pre-training phase}.

The upstream embeddings $e_U$ are removed and a new embedding layer, $e_D = (e_D^s, e_D^t)$ is initialized and prepended to $Y$. The weights of $Y$ are frozen, but $e_D$ is trained on $D_\text{train}$ in order to fit a lexical alignment to the representations learned during pre-training. This phase is called the \emph{fine-tuning phase}. During this phase, only a linear embedding layer is learned so any higher order features must come from structural representations learned upstream. Freezing $Y$ prevents leakage from the downstream task that would dilute the structural content learned during pre-training.

Finally, the composite model $Y\circ e_D$ is paired with a (fixed) decoding algorithm and evaluated on $D_\text{test}$. This is called the \emph{evaluation phase}.

\subsection{Evaluation}

We evaluate three performance metrics: (1) an approximation of semantic accuracy from \citet{nlc2cmd}; this is a score between --100 (worst) and 100 (best) which we call the Bash Similarity Heuristic (BaSH); (2) accuracy, which is the proportion of words which are the model's maximum likelihood estimate; (3) perplexity ($\ppl$) computed on downstream target programs.

\subsection{Datasets}

\paragraph{Copy task.}
$\sim$10k synthetic source sequences paired with their image under a vocabulary permutation.

\paragraph{End-to-End.}
NLBash \citep{nlc2cmd}. This is equivalent to end-to-end training.

\paragraph{English to Python.}
The CoNaLa bitext corpus from \citet{conala}.

\paragraph{English to SQL.}
The Spider dataset \citep{spider}. We mask constants and omit schema context to reduce complexity without disturbing syntax.

\paragraph{SQL to English (SQLNL).}
NLSQL with sources and targets interchanged.

\paragraph{English to German.}
The WMT14 English to German translation task (En-De) as curated by \citet{luong-etal-2015-effective}.

\paragraph{Reversal.}
$\sim$10k synthetic source sequences paired with targets constructed by reversing word order in the source.

\section{Results and Discussion}

\begin{table}[h]
\centering
\begin{tabular}{lrrr}
\toprule
\textbf{Experiment} & \textbf{ppl} & \textbf{Acc.} & \textbf{BaSH}\\ \hline
Uniform Init. & 3.29 & 67.9 & --8.3  \\
Xavier Init. & 3.11 & 75.0 & 15.1 \\
End-to-End & 3.80 & 78.1 & 16.5 \\ \hline 
NLPython & 3.39 & 74.2 & 10.1 \\
NLSQL & 3.14 & 75.6 & {\bf 14.1} \\
SQLNL & 3.27 & 75.2 & 13.5 \\
En-De & {\bf 2.77} & {\bf 76.3} & 12.9 \\
Reversal & 3.33 & 73.2 & 7.2 \\
Copy & 3.24 & 73.8 & 11.1 \\ \bottomrule
\end{tabular}
\caption{Results.}
\label{tab:results}
\end{table}

We find that representations from NLSQL transfer strongly to NLBash as is shown in Table \ref{tab:results}. This corroborates the intuition that these tasks are structurally similar and is expected since NLBash consists largely of \verb|find| commands which have a similar structure to \verb|SELECT| statements in SQL.

Even so, the model pre-trained on NLSQL did not score better than a model with random weights. This reinforces the somewhat surprising effectiveness of random models under the Xavier initialization scheme.

The copy task also scores well despite its simplicity. Indeed, we observe no substantial improvement of NLPython over the copy task, which suggests that \emph{although NLPython exhibits semantic patterns such as loops and control flow, it provides little transferable gain over direct lexical mapping}.

Finally, $\ppl$ and accuracy only seem loosely related to semantic accuracy measured by BaSH.
En-De outperformed all upstream tasks on $\ppl$ and accuracy, however, considerably more compute was expended during pre-training. This indicates a correlation between expended compute and perplexity, and the remaining measurements suggest this trend may not be shared by semantic accuracy. To test this idea, we varied dataset size and training length.

\subsection{Varying Compute}\label{sec:compute}

\begin{table}[h]
\centering
\begin{tabular}{lcccr}
\toprule
\textbf{Samples} & \textbf{Steps} & \textbf{ppl} & \textbf{Acc.} & \textbf{BaSH} \\ \hline
10k & 2500 & 3.39 & 75.0 & 13.0 \\
10k & 7500 & 3.27 & 75.0 & 11.8 \\
100k & 2500 & 3.13 & 76.0 & \textbf{14.2} \\
100k & 7500 & 3.16 & 75.8 & 13.7 \\
4.4M & 2500 & 2.95 & 76.0 & 12.6 \\
4.4M & 10000 & \textbf{2.77} & \textbf{76.3} & 12.9 \\
\bottomrule
\end{tabular}
\caption{English to German, several compute samples.}
\label{tab:decompute}
\end{table}

We randomly downsample En-De to 10K and 100K samples and train for 2500 and 7500 steps and display results in Table \ref{tab:decompute}.

There is a clear correlation between larger datasets, more training steps, and better performance measured by perplexity and accuracy. 
Interestingly, the same trend is not shared by BaSH. We instead find that \emph{transferability of semantic representations actually appears to degrade with upstream training past a point}.

\subsection{Future Work}

A natural extensions is to include (more) multilingual and multimodal upstream data.
Further, the extent that transferability matures into performance of fine-tuned models is untested. It would be valuable to investigate how the convergence rate of unfrozen, fine-tuned models varies with (1) upstream task and (2) compute expended during pre-training.

Finally, it would be valuable for practitioners to understand the strength of the correlation between transferability (measured here by BaSH evaluated on frozen transformers) and fine-tuned performance.

% ==============================================================================
%   End Content
% ==============================================================================

% Entries for the entire Anthology, followed by custom entries
\bibliography{anthology,custom}

\appendix

\section{Model}
\label{sec:model}

We use a simple transformer encoder decoder stack \citep{vaswani2017} with 6 layers and 8 attention heads, trained with cross-entropy loss. The hidden size is 512 and feedforward layers have dimension 2048. All models are regularized with dropout and label smoothing rates of 0.1. All models are trained on two NVIDIA GeForce GTX 1080 Ti GPUs with 11GB of random access memory and converge well within one day of training.

\end{document}